# Complexity transitions in chaotic quantum systems

### v1

@@ I Introduction
The interplay between interactions and disorder Vojta (2019) in quantum many-body systems gives rise to complex systems with extremely rich phenomenology, characterized by varying degrees of chaoticity and ergodicity Gornyi and Mirlin (2002); Pino et al. (2016); Basko et al. (2006); Pino et al. (2016); Green (2015); Abanin et al. (2019). Although analytical approaches are available in certain limits of ergodic Berry (1977); Bohigas et al. (1984); Haake (1991); Lahoche and Ousmane Samary (2020) and localized regimes Evers and Mirlin (2008); Hauke and Heyl (2015); Altman and Vosk (2015); Goremykina et al. (2019), and while quantum simulations Hauke et al. (2012); Georgescu et al. (2014); Daley et al. (2022); Fraxanet et al. (2023) have explored various regimes of disordered interacting systems Jurcevic et al. (2014); Schreiber et al. (2015); yoon Choi et al. (2016); Smith et al. (2016); Bordia et al. (2017); Lüschen et al. (2017); Rispoli et al. (2019); Sauerwein et al. (2023), their treatment remains a significant challenge both for theory and experiment. In order to translate that difficulty into a measurable quantity, various notions have been proposed. Notable examples of such “complexity markers” include entanglement entropy Horodecki et al. (2009) and stabilizer Rényi entropy Leone et al. (2022), which have clear implications both for the difficulty of numerically treating Vidal (2003); Eisert (2013); Orús (2014, 2019); Montangero et al. (2018); Mora and Briegel (2005); Bravyi et al. (2019); Chernyshev et al. (2024); Brökemeier et al. (2025) the system under study as well as for the quantum resources Chitambar and Gour (2019) an experiment needs in order to generate the target state Joshi et al. (2023); Howard and Campbell (2017). Such markers are also deeply connected to the question of what actually defines a quantum system as complex Liu et al. (2018); Baiguera et al. (2024); Bu et al. (2024); Haug et al. (2024); Leone et al. (2023). As a disordered quantum many-body system enters into respectively less ergodic and more localized regimes, the number of basis states involved is reduced (see Fig. 1b), suggesting intuitively a suppression of quantum complexity. As of yet, it remains, however, an open question how precisely ergodicity transitions are related to changes in complexity markers, and in particular, whether these define qualitatively different complexity regimes.
@@
In this paper, we map out complexity quantifiers (see Fig. 1) in prototypical disordered models that exhibit multiple distinct regimes: the Rosenzweig–Porter (RP) model Rosenzweig and Porter (1960); Altland et al. (1997), the Power-Law Random Banded Matrix (PLRBM) model Mirlin et al. (1996), and the Sachdev–Ye–Kitaev (SYK) Sachdev and Ye (1993); Kitaev (2015) model coupled to an Ising chain (see Fig. 2). The RP and PLRBM are paradigmatic models of chaotic systems that interpolate between ergodic and non-ergodic regimes, and they are central to the understanding of Anderson localization Anderson (1958); Evers and Mirlin (2000) and (multi)fractal states Evers and Mirlin (2008); Kravtsov et al. (2015); Mirlin and Evers (2000). The SYK model is a prototypical example of a maximally chaotic quantum many-body system Maldacena et al. (2016a), which plays a key role in the phenomenological description of non-Fermi liquid behavior Sachdev and Ye (1993); Chowdhury et al. (2022) and in the study of quantum black holes through the holographic principle Sachdev (2015); Cotler et al. (2017). We scrutinize these models using three different markers of complexity: (i) The fractal dimension Castellani and Peliti (1986); Halsey et al. (1986); Aoki (1986); Kramer and MacKinnon (1993); Čadež et al. (2024), a key observable for identifying localized regions and mobility edges, quantifies how much a state is spread over a computational basis. (ii) The entanglement entropy across half–half bipartitions estimates a state’s quantumness in terms of its inability to be expressed as a direct product state Orús (2019). However, not all entangled states are computationally challenging. An outstanding example is the class of stabilizer states, which contain highly-entangled states but can be tackled efficiently classically thanks to the Gottesman–Knill theorem Gottesman (1998a, b). In contrast, non-stabilizer states fall outside this framework and are exponentially hard to represent on a classical computer. To capture this distinction, we use as the third measure (iii) the non-stabilizerness Veitch et al. (2014) (or “magic”) of a state, quantified through the stabilizer Rényi entropy (SRE) Leone et al. (2022). While the fractal dimension and entanglement entropy have been instrumental in characterizing non-ergodic quantum systems already for some time Mirlin et al. (1996); Refael and Moore (2009); Lukin et al. (2019); Sierant et al. (2025), magic is currently emerging as a powerful tool in various contexts, including the study of ergodicity Vairogs and Yan (2024); Chen et al. (2024a); Zhang and Gu (2024); Turkeshi et al. (2023, 2025a, 2025b); Falcão et al. (2025), quantum chaos Leone et al. (2021); Goto et al. (2022); Varikuti et al. (2025), and as a sensitive probe for identifying quantum phase transitions White et al. (2021); Oliviero et al. (2022); Tarabunga et al. (2023); Falcão et al. (2024); Jasser et al. (2025); Ding et al. (2025); Hoshino et al. (2025).
@@
Finally, our studies complement ongoing research on several complexity markers in ergodic and integrable/localized systems, including Krylov complexity Rabinovici et al. (2022); Basu et al. (2024); Cohen et al. (2024); Sahu et al. (2024), logarithmic multifractality Chen et al. (2024b), operator spreading Nahum et al. (2018), holographic complexity Alishahiha (2015), or quantum-computational complexity Watrous (2008).
@@
Figure 1: Complexity markers across a transition from a chaotic (ergodic) to an integrable (localized) phase. (a) As the control parameter $\mu$ drives a random matrix model towards localization, the eigenvalue distribution evolves from the Wigner–Dyson semicircle to a Gaussian profile. (b) Comparison of three complexity markers (data for the RP model). A high (low) fractal dimension $D_{2}$ indicates delocalization (localization) in the computational basis. High (low) magic $\mathcal{M}_{2}$ implies support over an extensive set of $4^{N}$ Pauli strings (concentration on a limited subset). Strong (weak) entanglement $S_{\mathrm{vN}}$ reflects high (low) Schmidt rank and the presence of many (few) computational basis states in the superposition. While all three markers feature a high-complexity regime, the transition towards intermediate- or low-complexity regimes can happen at distinct points ($\mu_{1,2}^{D_{2}}\neq\mu_{1,2}^{\mathcal{M}_{2}}\neq\mu_{1,2}^{S}$). Regimes of different complexity thus overlap but do not coincide. (c) The phase transition points in the RP model, obtained through finite-size scaling analysis, differ across the three complexity markers (orange: high [$\mathcal{O}(1)$] complexity; green: intermediate; blue: low complexity [$\mathcal{O}(1/N)$]). Faded regions indicate numerical uncertainty in determining regime boundaries. Notably, the ground state exhibits a sharp transition directly from high to low complexity, bypassing the intermediate regime. In contrast, bulk eigenstates display a well-defined intermediate phase, with a significant variation in transition points across the considered markers.
@@ II.1 Fractal dimension
The fractal dimension can be derived from the inverse participation ratio (IPR), and quantifies the spread of the wavefunction in a given basis, which in our case will be given by the computational basis states {|$\{\ket{i}\}^{L}_{i=1}$, $L$ being the Hilbert space dimension. Fractal dimension and IPR are commonly used markers to investigate many-body localization Abanin et al. (2019); Castellani and Peliti (1986); Mirlin and Evers (2000); Evers and Mirlin (2000); Čadež et al. (2024), as a state that has support on just a few elements of the basis will be in the localized phase, while an ergodic wavefunction will be spread over many states of any generic basis, as shown in Fig. 1b. In computational terms, a highly localized state implies a smaller part of the Hilbert space needs to be handled, making it easier to simulate classically Prelovšek et al. (2018); Pietracaprina and Laflorencie (2021).
For a wavefunction |$\ket{\psi}$ expressed in terms of the computational basis with coefficients $\psi(i)=\braket{i}{\psi}$, the $q$th IPR is defined as
@@ II.2 Entanglement entropy
Given a bipartition of the Hilbert space $\mathcal{H}=\mathcal{H}_{A}\otimes\mathcal{H}_{B}$ and a pure state |$\ket{\psi}\in\mathcal{H}$, the von Neumann entropy is defined as $S(\rho_{A})=-\mathrm{Tr}(\rho_{A}\log\rho_{A})$, with $\rho_{A}=\mathrm{Tr}_{B}\ket{\psi}\bra{\psi}$ | the reduced density matrix for subsystem $A$. The maximal entanglement entropy is then defined as the von Neumann entropy maximized over all possible half–half bipartitions of the system,
@@ II.3 Stabilizer Rényi entropy (SRE)
The last complexity marker we consider is the stabilizer Rényi entropy (SRE). The SRE captures how a state is distributed when expressed in the basis of Pauli operators $P\in\mathcal{P}_{N}$, where $\mathcal{P}_{N}$ represents the Pauli group for $n$-qubits. Stabilizer states, which are common eigenstates of a maximal set of mutually commuting Pauli operators, exhibit a highly concentrated probability distribution of the Pauli expectation values, confined to a subset of $2^{N}$ operators, and thus are easily simulatable Gottesman (1997, 1998a); Aaronson and Gottesman (2004); Nest (2008); Gidney (2021); Bravyi et al. (2019). In contrast, a Haar-random state has approximately equal weight on all $P$’s Turkeshi et al. (2025a). Thus, a natural way to quantify the nonstabilizerness of a state |$\ket{\psi}$ is by assessing the spread of this distribution using the $q$-th order nonstabilizer Rényi entropy,
@@ III.1 Model
The relative simplicity of the RP model enables an analytical treatment of both its spectral properties  Kunz and Shapiro (1998); Kravtsov et al. (2015) and eigenvector statistics  von Soosten and Warzel (2019); Bogomolny and Sieber (2018a); Truong and Ossipov (2016), allowing for a comprehensive characterization of the model in the bulk of the energy spectrum. In the ergodic phase, which spans the range $10\leq\gamma<$ 1, the model behaves like an RMT, with Wigner–Dyson level spacing and Porter–Thomas statistics for the eigenvector probability distribution, signaling quantum chaos and ergodicity in Hilbert space. At $\gamma=1$, a transition to non-ergodic behavior occurs, and at $\gamma=2$ Anderson localization sets in. In the regime $2<\gamma<\infty$, wavefunctions are localized within a small region of Hilbert space, and the level spacing distribution exhibits Poissonian statistics. A key feature of the RP model is the presence of an intermediate regime between the ergodic and localized phases, in the range $21<\gamma<$ 2, where wavefunctions remain extended but occupy only a fraction of the full Hilbert space, exhibiting fractal behavior that deviates from conventional ergodicity without being fully localized. The fractal dimension defined in Eq. (3) is a particularly effective marker for these transitions Kravtsov et al. (2015), leading to the characterization
@@
Figure 3: (a1-c1) Complexity markers (normalized with respect to $N$), averaged over the central $20\%$ of the bulk eigenstates of the RP model for different system sizes from $N=4$ up to $N=12$, and averaged over samples from $20000$ for $N=4$ to $500$ for $N=12$. All markers plateau at the maximum value in an extended regime at small $\gamma$, overlapping with the model’s ergodic regime. The $N$-dependent value matches with predictions from the GUE ensemble that is exactly recovered at $\gamma=0$. The plateau’s extent in terms of $\gamma$ depends on the marker: the plateau is largest for $S_{\rm vN}$ and smallest for $D_{2}$. At large $\gamma$, the model enters a localized regime, characterized by a low value of all markers. The slope at which the model goes from the ergodic to the localized regime is different for all markers, as becomes clear from the derivatives (a2-c2). The crossings in the derivatives (insets in a2-c2) determine two different transition points $\gamma_{1}$ and $\gamma_{2}$. There are, therefore, three different regimes of high, intermediate, and low complexity, which do not coincide for the different complexity markers used.
@@ III.2 Results
For the fractal dimension, two critical points are known within $\gamma\in(0,3)$, corresponding to transitions from ergodic to non-ergodic at $\gamma_{1}^{c}=1$ and then to the localized phase at $\gamma_{2}^{c}=2$ [Eq. (10)]. However, it is not a priori clear that entanglement and magic should exhibit the identical behavior. As can be seen in Fig. 3b1,c1, at small $\gamma$ they exhibit $\mathcal{O}(1)$ scaling, indicating a high-complexity regime that one may expect in an ergodic phase. At a marker-dependent value $\gamma_{1}$, the complexity begins to decay, signaling entry into the non-ergodic regime. This decay continues until a second point $\gamma_{2}$, beyond which the markers eventually stabilize to a $\mathcal{O}(1/N)$ plateau as $N\to\infty$, indicating a low-complexity regime consistent with a large degree of localization. Within the intermediate regime $\gamma_{1}<\gamma<\gamma_{2}$, all markers display fractal-like scaling of the form $\mathcal{O}(N^{\delta})$ with $-1<\delta<0$.
@@ IV.2 Results
Figure 5: (a1-c1) Complexity markers (normalized with respect to $N$), averaged over the central $20\%$ of bulk eigenstates of the PLRBM model for $N=4$ up to $N=9$ ($N=12$ for the fractal dimension) averaged over different sample sizes of $35000$ to $3000$, respectively. Throughout an extended regime at small $\alpha$, all markers plateau at the value matching with predictions from the GUE ensemble (recovered by the model at $\alpha=0$). At large $\alpha$, the model transitions to a localized regime, characterized by a lower value of all the markers. The transition point can be deduced from the derivatives of the complexity markers (a2-c2). The crossings in the derivatives (insets a2-c2) effectively find only one distinct transition point. Therefore, in contrast to the RP model, there are only 2 different regimes of high and low complexity that coincide with the ergodic and localized phases, respectively.
@@
Although the analysis of the PLRBM model is restricted to smaller system sizes and has more sizable fluctuations, it supports what we conjectured for the RP model: the complexity markers assume distinct values across the ergodic-to-non-ergodic transition only in the presence of a genuine fractal (or multifractal) phase. Indeed, within error bars, in the PLRBM model, the transition points identified by all three complexity markers coincide (or almost coincide) with the Anderson critical point, the only transition captured by the fractal dimension in the thermodynamic limit. We do not report here the results for the ground state behavior of the PLRBM model since they qualitatively resemble that of the RP model considered in Sec. III.2: the transitions occur earlier than in the bulk, yet no intermediate phase is observed. This is consistent with expectations, as the markers used do not reveal a clear extended non-ergodic regime even in the bulk of the PLRBM model.

### v2

@@ I Introduction
The interplay between interactions and disorder Vojta (2019) in quantum many-body systems gives rise to complex systems with extremely rich phenomenology, characterized by varying degrees of chaoticity and ergodicity Gornyi and Mirlin (2002); Pino et al. (2016); Basko et al. (2006); Pino et al. (2016); Green (2015); Abanin et al. (2019). Although analytical approaches are available in certain limits of ergodic Berry (1977); Bohigas et al. (1984); Haake (1991); Lahoche and Ousmane Samary (2020) and localized regimes Evers and Mirlin (2008); Hauke and Heyl (2015); Altman and Vosk (2015); Goremykina et al. (2019), and while quantum simulations Hauke et al. (2012); Georgescu et al. (2014); Daley et al. (2022); Fraxanet et al. (2023) have explored various regimes of disordered interacting systems Jurcevic et al. (2014); Schreiber et al. (2015); yoon Choi et al. (2016); Smith et al. (2016); Bordia et al. (2017); Lüschen et al. (2017); Rispoli et al. (2019); Sauerwein et al. (2023), their treatment remains a significant challenge both for theory and experiment. In order to translate that difficulty into a measurable quantity, various notions have been proposed Osborne (2012); Orús (2019); Baiguera et al. (2025); Huang et al. (2024). Notable examples of such “complexity markers” include entanglement entropy Horodecki et al. (2009) and stabilizer Rényi entropy Leone et al. (2022), which have clear implications both for the difficulty of numerically treating Vidal (2003); Eisert (2013); Orús (2014, 2019); Montangero et al. (2018); Mora and Briegel (2005); Bravyi et al. (2019); Chernyshev et al. (2024); Brökemeier et al. (2025) the system under study as well as for the quantum resources Chitambar and Gour (2019) an experiment needs in order to generate the target state Joshi et al. (2023); Howard and Campbell (2017). Such markers are also deeply connected to the question of what actually defines a quantum system as complex Liu et al. (2018); Baiguera et al. (2024); Bu et al. (2024); Haug et al. (2024); Leone et al. (2023). As a disordered quantum many-body system enters into respectively less ergodic and more localized regimes, the number of basis states involved is reduced (see Fig. 1b), suggesting intuitively a suppression of quantum complexity. As of yet, it remains, however, an open question how precisely ergodicity transitions are related to changes in complexity markers, and in particular, whether these define qualitatively different complexity regimes.
@@
In this paper, we map out complexity quantifiers (see Fig. 1) in prototypical disordered models that exhibit multiple distinct regimes: the Rosenzweig–Porter (RP) model Rosenzweig and Porter (1960); Altland et al. (1997), the Power-Law Random Banded Matrix (PLRBM) model Mirlin et al. (1996), and the Sachdev–Ye–Kitaev (SYK) Sachdev and Ye (1993); Kitaev (2015) model coupled to an Ising chain (see Fig. 2). The RP and PLRBM are paradigmatic models of chaotic systems that interpolate between ergodic and non-ergodic regimes, and they are central to the understanding of Anderson localization Anderson (1958); Evers and Mirlin (2000) and (multi)fractal states Evers and Mirlin (2008); Kravtsov et al. (2015); Mirlin and Evers (2000). The SYK model is a prototypical example of a maximally chaotic quantum many-body system Maldacena et al. (2016a), which plays a key role in the phenomenological description of non-Fermi liquid behavior Sachdev and Ye (1993); Chowdhury et al. (2022) and in the study of quantum black holes through the holographic principle Sachdev (2015); Cotler et al. (2017). We scrutinize these models using three different markers of complexity: (i) The fractal dimension Castellani and Peliti (1986); Halsey et al. (1986); Aoki (1986); Kramer and MacKinnon (1993); Čadež et al. (2024), a key observable for identifying localized regions and mobility edges, quantifies how much a state is spread over a computational basis. (ii) The entanglement entropy across half–half bipartitions estimates a state’s quantumness in terms of its inability to be expressed as a direct product state Orús (2019). However, not all entangled states are computationally challenging. An outstanding example is the class of stabilizer states, which contain highly-entangled states but can be tackled efficiently classically thanks to the Gottesman–Knill theorem Gottesman (1998a, b). In contrast, non-stabilizer states fall outside this framework and are exponentially hard to represent on a classical computer. To capture this distinction, we use as the third measure (iii) the non-stabilizerness Veitch et al. (2014) (or “magic”) of a state, quantified through the stabilizer Rényi entropy (SRE) Leone et al. (2022). While the fractal dimension and entanglement entropy have been instrumental in characterizing non-ergodic quantum systems already for some time Mirlin et al. (1996); Refael and Moore (2009); Lukin et al. (2019); Sierant et al. (2025), magic is currently emerging as a powerful tool in various contexts, including the study of ergodicity Vairogs and Yan (2024); Chen et al. (2024a); Zhang and Gu (2024); Turkeshi et al. (2023, 2025a, 2025b); Falcão et al. (2025), quantum chaos Leone et al. (2021); Goto et al. (2022); Varikuti et al. (2025), and as a sensitive probe for identifying quantum phase transitions White et al. (2021); Oliviero et al. (2022); Tarabunga et al. (2023); Falcão et al. (2024); Jasser et al. (2025); Ding et al. (2025); Hoshino et al. (2025); Bera and Schirò (2025).
@@
Finally, our studies complement ongoing research on several complexity markers in ergodic and integrable/localized systems, including Krylov complexity Rabinovici et al. (2022); Basu et al. (2024); Cohen et al. (2024); Sahu et al. (2024); Bhattacharjee and Nandy (2025), logarithmic multifractality Chen et al. (2024b), operator spreading Nahum et al. (2018), holographic complexity Alishahiha (2015), or quantum-computational complexity Watrous (2008).
@@
Figure 1: Complexity markers across a transition from a chaotic (ergodic) to an integrable (localized) phase. (a) As the control parameter $\mu$ drives a random matrix model towards localization, the eigenvalue distribution evolves from the Wigner–Dyson semicircle to a Gaussian profile. (b) Comparison of three complexity markers (data for the RP model). A high (low) fractal dimension $D_{2}$ indicates delocalization (localization) in the computational basis. High (low) magic $\mathcal{M}_{2}$ implies support over an extensive set of $4^{N}$ Pauli strings (concentration on a limited subset). Strong (weak) entanglement $S_{\mathrm{vN}}$ reflects high (low) Schmidt rank and the presence of many (few) computational basis states in the superposition. While all three markers feature a high-complexity regime, the transition towards intermediate- or low-complexity regimes can happen at distinct points ($\mu_{1,2}^{D_{2}}\neq\mu_{1,2}^{\mathcal{M}_{2}}\neq\mu_{1,2}^{S}$). Regimes of different complexity thus overlap but do not coincide. (c) The phase transition points in the RP model, obtained through finite-size scaling analysis, differ across the three complexity markers (orange: high $\mathcal{O}(1)$] complexity; green: intermediate; blue: low complexity $\mathcal{O}(1/N)$]). Faded regions indicate numerical uncertainty in determining regime boundaries. Notably, the ground state exhibits a sharp transition directly from high to low complexity, bypassing the intermediate regime. In contrast, bulk eigenstates display a well-defined intermediate phase, with a significant variation in transition points across the considered markers.
@@ II.1 Fractal dimension
[…]
@@ II.2 Entanglement entropy
[…]
@@ II.3 Stabilizer Rényi entropy (SRE)
[…]
@@ III.1 Model
The relative simplicity of the RP model enables an analytical treatment of both its spectral properties  Kunz and Shapiro (1998); Kravtsov et al. (2015) and eigenvector statistics  von Soosten and Warzel (2019); Bogomolny and Sieber (2018a); Truong and Ossipov (2016), allowing for a comprehensive characterization of the model in the bulk of the energy spectrum. In the ergodic phase, which spans the range $0\leq\gamma<1$, the model behaves like an RMT, with Wigner–Dyson level spacing and Porter–Thomas statistics for the eigenvector probability distribution, signaling quantum chaos and ergodicity in Hilbert space. At $\gamma=1$, a transition to non-ergodic behavior occurs, and at $\gamma=2$ Anderson localization sets in. In the regime $2<\gamma<\infty$, wavefunctions are localized within a small region of Hilbert space, and the level spacing distribution exhibits Poissonian statistics. A key feature of the RP model is the presence of an intermediate regime between the ergodic and localized phases, in the range $1<\gamma<2$, where wavefunctions remain extended but occupy only a fraction of the full Hilbert space, exhibiting fractal behavior that deviates from conventional ergodicity without being fully localized. The fractal dimension defined in Eq. (3) is a particularly effective marker for these transitions Kravtsov et al. (2015), leading to the characterization
@@
Figure 3: (a1-c1) Complexity markers (normalized with respect to NN), averaged over the central $20\%$ of the bulk eigenstates of the RP model for different system sizes from $N=4$ up to $N=12$, and averaged over samples from $20000$ for $N=4$ to $500$ for $N=12$. All markers plateau at the maximum value in an extended regime at small $\gamma$, overlapping with the model’s ergodic regime. The NN-dependent value matches with predictions from the GUE ensemble that is exactly recovered at $\gamma=0$. The plateau’s extent in terms of $\gamma$ depends on the marker: the plateau is largest for $S_{\rm vN}$ and smallest for $D_{2}$. At large $\gamma$, the model enters a localized regime, characterized by a low value of all markers. The slope at which the model goes from the ergodic to the localized regime is different for all markers, as becomes clear from the derivatives (a2-c2). The crossings in the derivatives (insets in a2-c2) determine two different transition points $\gamma_{1}$ and $\gamma_{2}$. There are, therefore, three different regimes of high, intermediate, and low complexity, which do not coincide for the different complexity markers used.
@@ III.2 Results
For the fractal dimension, two critical points are known within $\gamma\in(0,3)$, corresponding to transitions from ergodic to non-ergodic at $\gamma_{1}^{c}=1$ and then to the localized phase at $\gamma_{2}^{c}=2$ [Eq. (10)]. However, it is not a priori clear that entanglement and magic should exhibit the identical behavior. As can be seen in Fig. 3b1,c1, at small $\gamma$ they exhibit $\mathcal{O}(1)$ scaling, indicating a high-complexity regime that one may expect in an ergodic phase. At a marker-dependent value $\gamma_{1}$, the complexity begins to decay, signaling entry into the non-ergodic regime. This decay continues until a second point $\gamma_{2}$, beyond which the markers eventually stabilize to a $\mathcal{O}(1/N)$ plateau as $N\to\infty$, indicating a low-complexity regime consistent with a large degree of localization. Within the intermediate regime $\gamma_{1}<\gamma<\gamma_{2}$, all markers display fractal-like scaling of the form $\mathcal{O}(N^{\delta})$ with $1<\delta<0$.
@@ IV.2 Results
Figure 5: (a1-c1) Complexity markers (normalized with respect to NN), averaged over the central $20\%$ of bulk eigenstates of the PLRBM model for $N=4$ up to $N=9$ ($N=12$ for the fractal dimension) averaged over different sample sizes of $35000$ to $3000$, respectively. Throughout an extended regime at small $\alpha$, all markers plateau at the value matching with predictions from the GUE ensemble (recovered by the model at $\alpha=0$). At large $\alpha$, the model transitions to a localized regime, characterized by a lower value of all the markers. The transition point can be deduced from the derivatives of the complexity markers (a2-c2). The crossings in the derivatives (insets a2-c2) effectively find only one distinct transition point. Therefore, in contrast to the RP model, there are only 2 different regimes of high and low complexity that coincide with the ergodic and localized phases, respectively.
@@
Although the analysis of the PLRBM model is restricted to smaller system sizes and has more sizable fluctuations, it supports what we conjectured for the RP model: the complexity markers assume distinct values across the ergodic-to-non-ergodic transition only in the presence of a genuine fractal (or multifractal) phase. Indeed, within error bars, in the PLRBM model, the transition points identified by all three complexity markers coincide (or almost coincide) with the Anderson critical point, the only transition captured by the fractal dimension in the thermodynamic limit. We do not report here the results for the ground state behavior of the PLRBM model since they qualitatively resemble that of the RP model considered in Sec. III.2: the transitions occur earlier than in the bulk Buijsman et al. (2025), yet no intermediate phase is observed. This is consistent with expectations, as the markers used do not reveal a clear extended non-ergodic regime even in the bulk of the PLRBM model.